OpenStudy (ariellesircy):

Multiply Rational Expressions I have my notes the teacher gave me but when I attempt to do these problems I'm doin something wrong. I'm supposed to find what x cannot equal. 21x-9/15x^2 • x^2-x-56/3x+21

OpenStudy (loser66):

\[\dfrac{21x -9}{x^2}*\dfrac{x^2-x-56}{3x+21}\] Right?

OpenStudy (loser66):

Hey, don't throw it here and let others do your homework.

OpenStudy (ariellesircy):

I didn't dump it here for you to do. I was hoping that someone would do it step by step and explain it so I could do my other ones.

OpenStudy (ariellesircy):

Can you please help me with this? @loser66

zepdrix (zepdrix):

Hey Arielle :) \[\Large\rm \frac{21x -9}{x^2}\cdot\frac{x^2-x-56}{3x+21}\]We need to find restrictions on x?

OpenStudy (ariellesircy):

Yes... like what x cannot equal.

zepdrix (zepdrix):

In the land of math, there are some really dirty dingey places that you want to avoid...\[\Large\rm \sqrt{-4} \quad\leftarrow\quad the~ghetto\]bad bad place, don't go there.

OpenStudy (ariellesircy):

the 21x-9 should be over 15x^2 -35x^3 I guess i must have messed that up when i typed it

zepdrix (zepdrix):

\[\Large\rm \frac{4}{0}\quad\leftarrow\quad docks~at~night\]Another bad place. Don't go here. We can't divide by 0 in the land of math, ya? That's what leading to restrictions in our problem. Certain values of x will lead to a 0 showing up in the bottom.

zepdrix (zepdrix):

\[\Large\rm \frac{12x-9}{15x^2 -35x^3}\cdot\frac{x^2-x-56}{3x+21}\]Like this? :) Oh ok

OpenStudy (ariellesircy):

Yes :)

OpenStudy (ariellesircy):

If you coukd explain what youre doing and why thatd be great!!!

zepdrix (zepdrix):

We have a problem when our equation looks like this:\[\Large\rm \frac{12x-9}{0}\cdot\frac{x^2-x-56}{3x+21}\]Or when our equation looks like this:\[\Large\rm \frac{12x-9}{15x^2 -35x^3}\cdot\frac{x^2-x-56}{0}\]Zeros in the bottom, bad bad bad. So this is a little nerve wracking i know :O We're like the police, we're going to go to that scary place... and then we'll say, "no! you can't do that!"

zepdrix (zepdrix):

So we'll let one of our factors in the bottom equal 0, \(\Large\rm 15x^2-35x^3=0\) and now we want to solve for x, we want to figure out which x values make this happen.

zepdrix (zepdrix):

So to solve for x here.. ummm looks like we need to do some factoring... each term has something in common. What can we factor out of them?

zepdrix (zepdrix):

What do you think? :) Gimme

OpenStudy (ariellesircy):

15x?

OpenStudy (ariellesircy):

15x^2?

zepdrix (zepdrix):

Mmm that's pretty close! :) Yes, we can pull 2 of the x's out. But we can't take a 15 out of 35, I guess we'll have to settle with a 5, ya?

OpenStudy (ariellesircy):

that makes sense

zepdrix (zepdrix):

\[\Large\rm 15x^2-35x^3=0\]\[\Large\rm 5x^2(3-7x)=0\]Do you understand how to solve for x in this setup? Do you remember your Zero Factor Property?

OpenStudy (ariellesircy):

I have to move my 3 and -7x over, square both sides, then move my 7x back to combine like terms and then divide that by both sides?

zepdrix (zepdrix):

No no, too fancy. Zero Factor Property tells us that when we multiply two things together and get 0, at least one of them is zero. \[\Large\rm a\cdot b=0, \qquad\to\qquad a=0,\quad\text{and/or}\quad b=0\]When we deal with x like this, we say that both will be zero at different times. So from here,\[\Large\rm 5x^2(3-7x)=0\]We can jump to here,\[\Large\rm 5x^2=0\qquad\text{and}\qquad (3-7x)=0\]

zepdrix (zepdrix):

And we'll look for solutions separately. Process look familiar I hope? XD

OpenStudy (ariellesircy):

Yes i remember that

zepdrix (zepdrix):

5x^2=0, divide by 5, x^2=0 square root, x=0 Ok that gives us a solution, ya? How bout the other equation. Can you find the x value? :)

OpenStudy (ariellesircy):

Yes. for the other one I subtracted my 3 and divided by -7 and got 3/7

zepdrix (zepdrix):

Ok good good good. So we've found two values for x that are bad! We went to that bad place, and found em. So we can confidently say that \(\Large\rm x\ne0,\quad x\ne\frac{3}{7}\).

zepdrix (zepdrix):

We've found two restrictions for x. Our other denominator will give us the other ones.

OpenStudy (ariellesircy):

okayy

zepdrix (zepdrix):

\[\Large\rm 3x+21=0\]So repeat the process with the other denominator. Turns out this one is much simpler though hehe

OpenStudy (ariellesircy):

Factor? 3(x+7)

OpenStudy (ariellesircy):

and then equal it to zero and solve?

zepdrix (zepdrix):

You could, yes. But notice that both terms don't have x's in them. So we don't have to be as picky with this one. We could simply subtract 21 to the other side, divide by 3, ya?

OpenStudy (ariellesircy):

Yes so the answer i s-7?

zepdrix (zepdrix):

Ok great. So our restrictions will be:\[\Large\rm x\ne0,\quad x\ne\frac{3}{7},\quad x\ne-7\]

zepdrix (zepdrix):

Those are all different values that would cause a 0 to show up in the denominator.

zepdrix (zepdrix):

Make a little more sense I hope? 0_o Math is weird, i know i know :D

OpenStudy (ariellesircy):

Okayy. So my teacher had taught us a way to just factor and then cross out like terms and x could not equal the opposite of those... So if i factor out each part of the equation and do it that way would this be correct for an answer left from the euation? 3/5x^2 times x-8/3 with x not equaling 0, 3/7, -7

zepdrix (zepdrix):

So your instructions are to, not only find restrictions, but to also simplify? Ok then ya I guess you have to factor out the top and bottom and do a bunch of fun stuff. We already factored the bottoms earlier, so our equation is:\[\Large\rm \frac{12x-9}{5x^2(3 -7x)}\cdot\frac{x^2-x-56}{3(x+7)}\]Yah you would cross out like terms :) So ummm the top...

OpenStudy (ariellesircy):

well she had said factor those too and cross out each two terms alike so i crossed out x+7 and the 3-7x on both leaving 3/5x^2 times x-8/3

zepdrix (zepdrix):

No no no silly. You cross out something from the top, with something from the bottom. Those are both in the bottom, they don't cross out. And further more, they're not the same, so they wouldn't cross out anyway. I don't understand where this x-8/3 is coming from 0_o

zepdrix (zepdrix):

Oh I think I see what you're saying maybe.. ummm

zepdrix (zepdrix):

Oh it was a 21 lolol my bad

OpenStudy (ariellesircy):

well when i factored i got 3(7x-3) and if you solve that and the bottom denominator the equal the same so i can cross them out...

zepdrix (zepdrix):

\[\Large\rm \frac{21x-9}{5x^2(3 -7x)}\cdot\frac{x^2-x-56}{3(x+7)}\]Becomes,\[\Large\rm \frac{-3(3-7x)}{5x^2(3 -7x)}\cdot\frac{(x-8)(x+7)}{3(x+7)}\]You can't cross out a (7x-3) with a (3-7x). Do you understand why I factored a -3 out instead of a 3 in the upper left? :)

OpenStudy (ariellesircy):

yes i see now.

OpenStudy (ariellesircy):

Could you walk me through another problem, but division this time?

zepdrix (zepdrix):

yaaa maybe :) open a new question though >.< this one is getting really laggy with all of the LaTeX

OpenStudy (ariellesircy):

okayy.